# CoFiI2P: Coarse-to-Fine Correspondences-Based Image-to-Point Cloud Registration

### v3

@@ III-C I2P Fine Matching
The first-stage matching at the coarse level constructs coarse super-point/super-pixel pairs $\tilde{\Omega}_{\mathcal{M}}$ but leads to poor registration accuracy. In order to obtain high-quality I2P correspondences, we generate fine correspondence based on the coarse matching results. During the decoder process, in each super-point/super-pixel correspondence $(\tilde{p}_{x},\tilde{i}_{y})$, the super-point $\tilde{p}_{X}$ reverse to local points group $G_{\tilde{p}_{x}}$ and the super-pixel $\tilde{i}_{y}$ reverse to local pixels patch $G_{\tilde{i}_{y}}$. Considering the uneven distribution of point cloud and computational efficiency, only the node points $p_{n}$ in local patch group $G_{\tilde{p}}$ are selected to establish correspondences. For each node point $p_{n}$, we select the pixel $i_{k}\in G_{\tilde{i}_{y}}$ that lies nearest in the feature space. Point-pixel pairs in each super-point-super-pixel pair are staked together as the dense corresponding pairs $\Omega_{\mathcal{M}}$. With point feature map $\mathbf{F}_{G_{\tilde{p}}}$ and pixel feature map $\mathbf{F}_{G_{\tilde{i}}}$ of local patch, the fine matching process is defined as:
@@ III-E Loss Function
In order to supervise the network simultaneously learning coarse level descriptors, fine level descriptors and in/beyond-frustum super-points classification, we introduce a joint loss function consisting of coarse level descriptor loss $\mathcal{L}_{coarse}$, fine level descriptor loss $\mathcal{L}_{fine}$ and classification loss $\mathcal{L}_{classify}$. The classification loss encourages the network to correctly label each super-point, and two descriptor losses pull positive matching pairs closer and push negative matching pairs farther in feature space.
The cosine similarity rule $s(p_{x},i_{y})$ of point cloud feature vector $\mathbf{F}({p_{x}})$ and image feature vector $\mathbf{F}(i_{y})$ is defined as :
$s(p_{x},i_{y})=\frac{<\mathbf{F}({p_{x}}),\mathbf{F}(i_{y})>}{\|\mathbf{F}_{(}{p_{x}})\|\|\mathbf{F}(i_{y})\|},$ (7)
@@
where $\mathbf{T}$ is the ground-truth transform matrix from point cloud coordinate system to image frustum coordinate system, and $\Gamma$ represents the mapping function that converts points from camera frustum to image plane coordinate system. The negative anchor $\tilde{i}_{neg}$ is defined as the super-pixel with the smallest distance to the $\tilde{p}_{x}$ in the feature space:
$\tilde{i}_{neg}=\mathop{\mathrm{argmin}}\limits_{\tilde{i}\in\tilde{I}}{\|d(\tilde{p}_{x},\tilde{i}_{neg})\|}\quad s.t.\quad\|\tilde{i}_{neg}-\tilde{i}_{pos}\|>r,$ (10)
where $r$ is the safe radius to remove adversarial samples. Finally, with positive margin $\Delta_{pos}$ and negative margin $\Delta_{neg}$, coarse level descriptor loss is defined in a triplet way as Eq. (11) :
@@
$\begin{split}\mathcal{L}_{fine}=\log\bigg{(}&1+\exp\big{(}{-\gamma\alpha_{pos}(s(p_{x},i_{pos})-\Delta_{pos})}\big{)}\\ &\sum_{(p_{x},i_{pos},i_{neg})}\exp\big{(}{\gamma{\alpha}_{neg}(p_{x},s(i_{neg})-\Delta_{neg})}\big{)}\bigg{)},\end{split}$ (12)
@@
Super-points classification loss is a binary cross-entropy loss:
$\mathcal{L}_{classify}=-\sum_{\tilde{p}_{x}\in\tilde{P}}\big{(}\tilde{p}_{x}\log(\tilde{p}_{x}^{\star})+(1-\tilde{p}_{x})\log{(1-\tilde{p}_{x}^{\star})}\big{)}.$ (14)
Overall, our loss function is
@@ IV-A1 Dataset
we evaluate our method on KITTI Odometry dataset [35], a benchmark dataset extensively employed in the field of I2P registration. KITTI Odometry dataset consists of 22 sequences of images and point cloud data collected simultaneously in urban environments, encompassing a variety of road scenes and environmental conditions, and 11 sequences provide ground-truth calibration files. The camera intrinsic function $\Gamma$ extracted from calibration files is thought unbiased during experiments. For a fair comparison with existing approaches [7, 8], sequences 0-8 are used for training and 9-10 for testing.
@@ IV-A2 Baseline methods
we compare proposed CoFiI2P with two open-sourced I2P methods:
  • DeepI2P [7]: proposed the frustum classification and inverse camera projection to estimate the camera pose. Due to the time-consuming data pre-processing and training, we reference the registration results reported in the paper.
  • CorrI2P [8]: is SOTA comparable I2P method. CorrI2P [8] predicts the overlapping area and establishes correspondences densely for pose estimation. We use officially released codes to reimplement this method. The only difference is the number of points changed from 40960 to 20480 for fair comparison.
@@ IV-A3 Evaluation metrics
we report the relative rotation error (RRE), relative translation error (RTE) and registration recall (RR) to evaluate the registration results. Inlier ratio (IR) and root mean square error (RMSE) used in I2P and P2P approaches [22, 36] are introduced to evaluate the quality of correspondences. RRE and RTE are defined as :
@@
$\mathrm{RR}_{\mathcal{M}}=\frac{1}{\lvert\Omega_{\mathcal{M}}\rvert}\sum_{i=1}^{\lvert\Omega_{\mathcal{M}}\rvert}\mathbbm{1}(\sqrt{\frac{1}{\lvert\Omega_{\mathcal{M}}\rvert}\sum_{(p_{x},i_{y})\in\Omega_{\mathcal{M}}}\|\Gamma(\mathbf{T}p_{x})-i_{y}\|}<\tau),$ (17)
@@
$\mathrm{IR}_{\mathcal{M}}=\frac{1}{\lvert\Omega_{\mathcal{M}}\rvert}\sum_{(p_{x},i_{y})\in\Omega_{\mathcal{M}}}\mathbbm{1}(\|\Gamma(\mathbf{T}p_{x})-i_{y}\|<\tau),\vspace{-0.1cm}$ (18)
in which $\tau$ is used to control the reprojection error tolerance. RMSE denotes the average reprojection error over the correspondences set $\Omega_{\mathcal{M}}$:
$\mathrm{RMSE}_{\mathcal{M}}=\sqrt{\frac{1}{\lvert\Omega_{\mathcal{M}}\rvert}\sum_{(p_{x},i_{y})\in\Omega_{\mathcal{M}}}\|\Gamma(\mathbf{T}p_{x})-i_{y}\|}.$ (19)
IV-B Implementation Details
Raw images and point clouds are preprocessed to filter out noise and reduce the computational burden. The top 50 rows of images are cropped out as most of these pixels see the sky. Then images are resized to $512\times 160$ for training and testing. The points are downsampled with $0.1m\times 0.1m\times 0.1m$ voxel and then randomly sampled 20480 points as input.
During the training process, with correct transform parameters provided by the calibration files, the ground truth correspondences $\Omega_{\mathcal{M^{\star}}}$ are established to supervise the network. We trained the whole network 25 epochs with batch size of 1. We use the Adam to optimize the network, and the initial learning rate is 0.001 and multiple by 0.25 after every 5 epochs. For our joint loss, we set $\lambda_{1}=\lambda_{2}=\lambda_{3}=1$. The safe radius $r$, positive margin $\Delta_{pos}$ and negative margin $\Delta_{neg}$ in loss function are set to 1, 0.2 and 1.8 respectively. Scale factor $\gamma$ is set to 10. During the coarse matching, super-points are projected to the frustum of the image, and beyond-frustum super-points are removed. Then we uniformly sample 64 super-points from in-frustum ones. The corresponding super-pixels are found with Eq. 4. The training of CoFiI2P takes about 49 hours. During the testing process, we set the confidence threshold as 0.9 to obtain reliable in-frustum super-points in coarse matching. All the experiments are conducted on a single RTX 4090 GPU.
TABLE I: Registration accuracy on KITTI dataset
@@ IV-B Implementation Details
DeepI2P (2D) 10/5 $7.56\pm 7.63$ $3.28\pm 3.09$
DeepI2P (3D) 10/5 $15.52\pm 12.73$ $3.17\pm 3.22$
CorrI2P none/none $6.93\pm 29.03$ $2.68\pm 10.51$
45/10 $3.41\pm 3.64$ $1.48\pm 1.35$ 97.07
10/5 $2.70\pm 1.97$ $1.24\pm 0.87$ 90.66
CoFiI2P none/none $\boldsymbol{1.14}\pm\boldsymbol{0.78}$ $\boldsymbol{0.29}\pm\boldsymbol{0.19}$
45/10 $\boldsymbol{1.14}\pm\boldsymbol{0.78}$ $\boldsymbol{0.29}\pm\boldsymbol{0.19}$ 100.00
10/5 $\boldsymbol{1.14}\pm\boldsymbol{0.78}$ $\boldsymbol{0.29}\pm\boldsymbol{0.19}$ 100.00
IV-C Registration Accuracy
@@
We report the RRE and RTE of our CoFiI2P as registration evaluation metrics under three different settings in Table I, where $none/none$ means no specific thresholds for filtering out false registration frames and $10^{\circ}/5m$ and $45^{\circ}/10m$ mean that any frames with registration errors exceeding these specified thresholds are ignored during the evaluation process. By employing these diverse settings, we can conduct a more equitable analysis. As shown in Table I, the proposed method outperforms all baseline methods by a large margin. Specifically, with different registration error threshold $none/none,45^{\circ}/10m,10^{\circ}/5m$, the RRE reduces 84%, 67%, 58% and the RTE reduces 89%, 80%, 77% compared to CorrI2P [8] respectively. Meanwhile, the RR improves by 2.93%, 9.34% under two pairs of thresholds. Notably, our method achieves 100% RR even under hardest pair of thresholds $10^{\circ}/5m$, which indicates that our method achieves both robust and accurate performance in all the evaluation scenes. Furthermore, we project the points to the image plane with transform parameters provided by ground-truth files, CorrI2P [8] and CoFiI2P respectively and deploy the qualitative registration results in Fig. 5. It can be seen that the CoFiI2P remains stable and provides more accurate results, which confirms our claim.
Figure 3: Quantitative results of correspondences. (a) shows the inlier ratio, while (b) and (c) show the RMSE distribution of the CorrI2P and our CoFiI2P.
Figure 4: Qualitative results of correspondences. The first column shows the input, and the second and third column shows the correspondences of CorrI2P and our CoFiI2P respectively. The green lines represent correct matches, while the red lines represent incorrect matches.
@@
Figure 5: Quantitative registration results. The colors are rendered based on depth, ranging from blue in the foreground to red in the distance.
For the super-point frustum classification, the precision and recall of our method are 91.23% and 99.75% on the KITTI dataset, respectively. However, these numbers can not directly measure the quality of the correspondence. Therefore, we introduce the IR and RMSE metrics as in P2P registration approach [22] and show the results in Fig. 3 and Fig. 4. Fig. 3 (a) shows that CoFiI2P gets obviously higher IR scores than CorrI2P [8] under various thresholds, while Fig. 3 (b) and (c) demonstrate that our CoFiI2P owns higher correspondence accuracy. Specifically, CoFiI2P reduces the global average RMSE value from 30.35 pixels of CorrI2P [8] (seriously affected by the outliers with RMSE higher than 100 pixels) to 3.58 pixels. Fig. 4 shows that our CoFiI2P provides much cleaner correspondences set for pose estimation.
IV-D Ablation studies and Insights
@@
In this part, we analyze three crucial factors in our CoFiI2P: I2P transformer module, coarse-to-fine matching scheme and point cloud density. We conduct ablation studies on them to prove the effectiveness of each module and review the influence of point cloud density. In order to reduce the computational burden, we train the variant networks 10 epochs in the discussion part, instead of 25 epochs in the experiments part. We report the global RRE and RTE as evaluation metrics and no thresholds are used to reject false registration scenes.
IV-D1 Analyze of the I2P transformer

### v4

@@ III-C I2P Fine Matching
The first-stage matching at the coarse level constructs $robust$ super-point/super-pixel pairs $\tilde{\Omega}_{\mathcal{M}}$ but leads to poor registration accuracy. To obtain high-quality I2P correspondences, we generate fine correspondence based on the coarse matching results. In each super-point/super-pixel correspondence $(\tilde{p}_{x},\tilde{i}_{y})$, the super-point $\tilde{p}_{x}$ is expanded to a local points group $G_{\tilde{p}_{x}}$ and the super-pixel $\tilde{i}_{y}$ is expanded to a local pixels patch $G_{\tilde{i}_{y}}$. Considering the uneven geometric distribution of point cloud and computational efficiency, only the node points $p_{n}$ in local patch group $G_{\tilde{p}}$ are selected to establish correspondences. For each node point $p_{n}$, we select the pixel $i_{k}\in G_{\tilde{i}_{y}}$ that lies nearest in the feature space. Point-pixel pairs in each super-point-super-pixel pair are stacked together as the fine corresponding pairs $\Omega_{\mathcal{M}}$. With point feature map $\mathbf{F}_{G_{\tilde{p}}}$ and pixel feature map $\mathbf{F}_{G_{\tilde{i}}}$ of local patch, the fine matching process is defined as:
@@ III-E Loss Function
To learn the coarse level descriptors and fine level descriptors and in/beyond-frustum super-points classification simultaneously, we introduce a joint loss $\mathcal{L}$ consisting of coarse level descriptor loss $\mathcal{L}_{coarse}$ , fine level descriptor loss $\mathcal{L}_{fine}$ and classification loss $\mathcal{L}_{classify}$. The classification loss encourages the network to correctly label each super-point, and two descriptor losses pull positive matching pairs closer together and push negative matching pairs farther apart in the feature space.
The cosine similarity $s(p_{x},i_{y})$ of point cloud feature vector $\mathbf{F}_{p_{x}}$ and image feature vector $\mathbf{F}_{i_{y}}$ is defined as :
$s(p_{x},i_{y})=\frac{<{\color[rgb]{0,0,0}\mathbf{F}_{p_{x}}},{\color[rgb]{0,0,0}\mathbf{F}_{i_{y}}}>}{\|{\color[rgb]{0,0,0}\mathbf{F}_{p_{x}}}\|\|{\color[rgb]{0,0,0}\mathbf{F}_{i_{y}}}\|},$ (7)
@@
where $\mathbf{T}$ is the ground-truth transform matrix from point cloud coordinate system to image frustum coordinate system, and $\Gamma$ represents the mapping function that converts points from the camera frustum to image plane coordinate system. The negative anchor $\tilde{i}_{neg}$ is defined as the super-pixel with the smallest distance to the $\tilde{p}_{x}$ in the feature space:
$\tilde{i}_{neg}=\mathop{\mathrm{argmin}}\limits_{\tilde{i}\in\tilde{I}}{\|d(\mathbf{F}_{\tilde{p}_{x}},\mathbf{F}_{\tilde{i}_{neg}})\|}\quad s.t.\quad\|\tilde{i}_{neg}-\tilde{i}_{pos}\|>r,$ (10)
where $r$ is the safe radius to remove adversarial samples. Finally, with positive margin $\Delta_{pos}$ and negative margin $\Delta_{neg}$, coarse level descriptor loss is defined in a triplet way [27] as Eq. (11) :
@@
$\begin{split}\mathcal{L}_{fine}=\log\bigg{(}&1+\exp\big{(}{-\gamma\alpha_{pos}(s(p_{x},i_{pos})-\Delta_{pos})}\big{)}\\ &\sum_{(p_{x},i_{pos},i_{neg})}\exp\big{(}{\gamma{\alpha}_{neg}(p_{x},s({\color[rgb]{0,0,0}p_{x},i_{pos}})-\Delta_{neg})}\big{)}\bigg{)},\end{split}$ (12)
@@
Super-points classification is defined as a binary cross-entropy loss:
$\mathcal{L}_{classify}=-\sum_{\tilde{p}_{x}\in\tilde{P}}\big{(}{\color[rgb]{0,0,0}\hat{s}_{\tilde{p}_{x}}}\log({\color[rgb]{0,0,0}s_{\tilde{p}_{x}}})+(1-{\color[rgb]{0,0,0}\hat{s}_{\tilde{p}_{x}}})\log{(1-{\color[rgb]{0,0,0}s_{\tilde{p}_{x}}})}\big{)},$ (14)
To avoid ambiguity, we reuse $s_{\tilde{p}_{x}}$ for score prediction and $\hat{s}_{\tilde{p}_{x}}$ for ground truth label. Overall, our loss function is
@@ IV-A1 Dataset
We evaluate our method on two public datasets, KITTI Odometry [36] and Nuscenes [37].
  • KITTI Odometry [36]: It consists of 11 sequences of images and point cloud data in urban environments with ground-truth calibration files. The camera intrinsic function $\Gamma$ extracted from calibration files is thought unbiased during experiments. Sequences 0-8 are used for training and 9-10 for testing as previous approaches [7, 8]. The image resolution is resized to 160$\times$512 and the point cloud is downsampled and randomly selected to 20480.
  • Nuscenes [37]: The image and point cloud pairs are generated by the official SDK, where the point cloud is accumulated from the nearby frames and the image is from the current frame. We follow the official data split of nuScenes to utilize 850 scenes for training, and 150 scenes for testing. We downsample the image resolution to 160$\times$320 and the point cloud size to 20480 similarly.
@@ IV-A2 Baseline Methods
We compare the proposed CoFiI2P with two open-sourced I2P methods as follows.
  • DeepI2P [7]: It uses the frustum classification and inverse camera projection to estimate the camera pose. We use officially released code for reimplementation and use the 2D and 3D inverse camera projection for optimization, namely DeepI2P (2D) and DeepI2P (3D).
  • CorrI2P [8]: It is the SOTA I2P method. CorrI2P predicts the overlapping area and establishes correspondences densely for pose estimation. We use officially released code for reimplementation.
@@ IV-A3 Evaluation Metrics
We calculate relative rotation error (RRE), relative translation error (RTE) and registration recall (RR) to evaluate the registration accuracy. Inlier ratio (IR) used in previous approaches [21, 38] is introduced to evaluate the quality of correspondences. For efficiency analysis, we report the frame-per-second (FPS) evaluated on an Intel i9-13900K CPU and NVIDIA RTX 4090 GPU with batch size of 1. RRE and RTE are defined as :
@@
$\mathrm{RR}_{\mathcal{M}}=\frac{1}{\lvert\Omega_{\mathcal{M}}\rvert}\sum_{i=1}^{\lvert\Omega_{\mathcal{M}}\rvert}\mathds{1}(\sqrt{\frac{1}{\lvert\Omega_{\mathcal{M}}\rvert}\sum_{(p_{x},i_{y})\in\Omega_{\mathcal{M}}}\|\Gamma(\mathbf{T}p_{x})-i_{y}\|^{2}}<\tau),$ (17)
@@
$\mathrm{IR}_{\mathcal{M}}=\frac{1}{\lvert\Omega_{\mathcal{M}}\rvert}\sum_{(p_{x},i_{y})\in\Omega_{\mathcal{M}}}\mathds{1}(\|\Gamma(\mathbf{T}p_{x})-i_{y}\|<\tau),\vspace{-0.1cm}$ (18)
in which $\tau$ is used to control the reprojection error tolerance.
IV-A4 Implementation Details
With correct transform parameters provided by the calibration files during the training process, the ground truth correspondences $\Omega_{\mathcal{M^{\star}}}$ are established to supervise the network. We trained the whole network 25 epochs for the KITTI Odometry dataset and 10 epochs for the Nuscenes dataset. We use the Adam [39] to optimize the network, and the initial learning rate is 0.001 and multiple by 0.25 after every 5 epochs. For our joint loss, we set $\lambda_{1}=\lambda_{2}=\lambda_{3}=1$. The safe radius $r$, positive margin $\Delta_{pos}$ and negative margin $\Delta_{neg}$ in loss function are set to 1, 0.2 and 1.8. Scale factor $\gamma$ is set to 10. More details are available in open-source code.
TABLE I: Registration accuracy on KITTI Odometry [36] and Nuscenes [37] datasets. The results are presented in the format of "mean $\pm$ standard" deviation. $\uparrow$ means higher is better and $\downarrow$ means lower is better, respectively. The best results are highlighted in bold. Results of DeepI2P on the KITTI Odometry dataset are taken from [7] with identical settings.
@@ IV-A4 Implementation Details
RRE($\downarrow$ RTE(m)$\downarrow$ RR(%) $\uparrow$ RRE($\downarrow$ RTE(m)$\downarrow$ RR(%) $\uparrow$
DeepI2P (2D) [7] 10/5 $7.56\pm 7.63$ $3.28\pm 3.09$ $2.81\pm 2.23$ $1.48\pm 0.88$ 95.13 1.41
DeepI2P (3D) [7] 10/5 $15.52\pm 12.73$ $3.17\pm 3.22$ $6.60\pm 19.33$ $1.31\pm 1.57$ 38.64 0.71
CorrI2P [8] none/none $6.93\pm 29.03$ $2.68\pm 10.51$ $8.47\pm 24.02$ $5.30\pm 8.10$ 7.95
45/10 $3.41\pm 3.64$ $1.48\pm 1.35$ 97.07 $5.42\pm 4.30$ $3.78\pm 2.24$ 91.07
10/5 $2.70\pm 1.97$ $1.24\pm 0.87$ 90.66 $3.90\pm 2.23$ $2.61\pm 1.23$ 61.67
CoFiI2P none/none $\boldsymbol{1.14}\pm\boldsymbol{0.78}$ $\boldsymbol{0.29}\pm\boldsymbol{0.19}$ $\boldsymbol{2.04}\pm\boldsymbol{5.54}$ $\boldsymbol{0.95}\pm\boldsymbol{5.04}$ 15.43
45/10 $\boldsymbol{1.14}\pm\boldsymbol{0.78}$ $\boldsymbol{0.29}\pm\boldsymbol{0.19}$ 100.00 $\boldsymbol{1.88}\pm\boldsymbol{1.77}$ $\boldsymbol{0.81}\pm\boldsymbol{0.59}$ 99.84
10/5 $\boldsymbol{1.14}\pm\boldsymbol{0.78}$ $\boldsymbol{0.29}\pm\boldsymbol{0.19}$ 100.00 $\boldsymbol{1.79}\pm\boldsymbol{1.22}$ $\boldsymbol{0.79}\pm\boldsymbol{0.53}$ 99.18
IV-B Registration Accuracy
@@
We report the RRE and RTE as evaluation metrics under three different settings in Table VII, where $none/none$ means no specific thresholds for filtering out false registration frames and $10^{\circ}/5m$ and $45^{\circ}/10m$ mean that any frames with registration errors exceeding these specified thresholds are ignored during the evaluation process. As shown in Table VII, the proposed method outperforms all baseline methods on the RRE and RTE registration metrics. Notably, our method achieves 100% RR and 99.18% RR under the hardest pair of thresholds $10^{\circ}/5m$ on the KITTI Odometry and Nuscenes datasets respectively, which indicates that our method constructs robust correspondences and achieves accurate registration performance in most of evaluation scenes. Besides, we project the points to the image plane with transform parameters provided by ground-truth files, CorrI2P [8] and CoFiI2P respectively and deploy the qualitative registration results in Fig. 3. It shows that the CoFiI2P remains stable and provides more accurate results, which confirms our claim. Although our method is slightly slower during the model inference step, the pose estimation is extremely fast due to the small quantity and high quality of matching pairs, thereby still maintaining online performance.
@@
Figure 3: Quantitative registration results on the KITTI Odometry dataset. The colors are rendered based on depth, ranging from blue in the foreground to red in the distance.
Figure 4: Quantitative results of correspondences. (a) and (b) shows the inlier ratio of our method (blue line) and CorrI2P (orange line) on KITTI Odometry and Nuscenes dataset respectively.
Figure 5: Qualitative results of correspondences on the KITTI Odometry [36] dataset. The first column shows the input, and the second and third column shows the correspondences of CorrI2P and our CoFiI2P respectively. The green lines represent correct matches, while the red lines represent incorrect matches.
For the super-point frustum classification, We introduce the IR metrics to evaluate the quality of the correspondence as in the P2P registration approach [21]. IR curves in Fig. 5 indicate that our method achieves a higher proportion of correct matches among the established correspondences, which leads to better registration results. Fig. 5 intuitively illustrates that our CoFiI2P provides a much cleaner correspondence set than baseline method.
IV-C Ablation Studies and Analysis
@@
In this part, we analyze three crucial factors in our CoFiI2P: I2P transformer module, coarse-to-fine matching scheme, and point cloud density. We conduct ablation studies on the KITTI Odometry [36] dataset to prove the effectiveness of each module and review the influence of point cloud density. We train the ablation models for 25 epochs as in the experimental part, and all other settings remain the same. We report global RRE and RTE as evaluation metrics and no thresholds are used to reject false registration scenes.
IV-C1 Analysis of the I2P Transformer